# Revisiting Multi-Agent Asynchronous Online Optimization with Delays: the Strongly Convex Case

### v1

@@ 1 Introduction
Motivated by these applications, Hsieh et al. (2022) extend the standard OCO into multi-agent asynchronous online optimization with delays. The setting involves a set of agents =$\mathcal{M}=\{1,\ldots,M\}$, and only one agent becomes active in each round $t$ to make the decision $\mathbf{x}_{t}$. Due to the effect of delays and asynchronicities, each agent $i\in\mathcal{M}$ can only receive the feedback about $f_{t}(\cdot)$ at the end of round $t+d_{t,i}-1$, where $d_{t,i}\in\mathbb{Z}^{+}$ denotes an arbitrary delay. It is well-known that if only one agent exists, i.e., $M=1$, various algorithms have been proposed to achieve an optimal $O(\sqrt{dT})$ regret bound for convex functions and an optimal $O(d\log T)$ regret bound for strongly convex functions in the worst case (Weinberger and Ordentlich, 2002; Joulani et al., 2013; Quanrud and Khashabi, 2015; Joulani et al., 2016; Wan et al., 2022b), where $d=\max_{t,i}\{d_{t,i}\}$ denotes the maximum delay. However, compared with the single-agent case, multi-agent OCO with delays faces two new challenges: the non-monotonicity of feedback sequence and lack of synchronization, which make the time-decaying and even delay-dependent learning rates utilized in these single-agent delayed algorithms unavailable (Hsieh et al., 2022).
@@ 2.1 The Standard OCO
Since the pioneering work of Zinkevich (2003), there has been a growing research interest in developing and analyzing algorithms for OCO (Hazan et al., 2007; Abernethy et al., 2008; Duchi et al., 2011; Kingma and Ba, 2015; Wang et al., 2020; Wan and Zhang, 2021, 2022). Roughly speaking, these existing OCO algorithms can be divided into two categories. The first category includes online gradient descent (OGD) (Zinkevich, 2003) and its variants, which can be viewed as an extension from gradient descent in offline optimization (Boyd and Vandenberghe, 2004) into the online setting. The other category contains follow-the-leader (FTL) (Hannan, 1957) and its variants, where the so-called “leader” is the optimal decision for all historical loss functions. Note that both categories are able to achieve an $O(\sqrt{T})$ regret bound for convex functions, and an $O(\log T)$ regret bound for strongly convex functions (Zinkevich, 2003; Hazan et al., 2007; Hazan, 2016). Moreover, these $O(\sqrt{T})$ and $O(\log T)$ regret bounds have been proved to be optimal for convex and strongly convex functions, respectively (Abernethy et al., 2008).
@@ 2.2 Multi-Agent OCO with Delays
where $\mathcal{F}_{t}$ is a set including the timestamps of all feedback received by the $t$-th active agent. They first prove that DDA can achieve the $O(\sqrt{dT})$ regret bound for convex functions, by setting
@@ 3.2 Follow-the-Delayed-Leader
Before introducing our algorithms, we first provide a formal definition for the available information held by each agent $i\in[M]$. To be precise, let $\mathcal{G}_{i}$ denote the set including the timestamps for all feedback received by each agent $i\in[M]$. In the beginning, the set should be initialized as $\mathcal{G}_{i}=\emptyset$. Then, due to the effect of delays and asynchronicity, each agent $i$ will receive the feedback from some previous round $s$, and update the set as $\mathcal{G}_{i}=\mathcal{G}_{i}\cup\{s\}$. Moreover, it will relay the received feedback if necessary. In this way, when an agent $i$ is activated at round $t$, it can only make the decision based on the feedback from rounds in the latest set $\mathcal{G}_{i}$ that may only be a subset of $[t-1]$.
Algorithm 1 Follow-the-Delayed-Leader (Agent $i$)
@@
3:     Asynchronously receive feedback $f_{s}(\mathbf{x})$ from round $s$
4:     $\mathcal{G}_{i}=\mathcal{G}_{i}\cup\{s\}$
5:     Relay $f_{s}(\mathbf{x})$ if necessary
6:     if the agent becomes active then
7:        $\mathcal{F}_{t}=\mathcal{G}_{i}$
8:        Play $\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}f_{s}(\mathbf{x})$
9:     end if
10:  end while
Therefore, even if the functions are strongly convex, FTL in (1) cannot be directly utilized. To address this issue, we propose a delayed variant of FTL for multi-agent OCO with delays that is called follow-the-delayed-leader (FTDL) and utilizes the following decision
where $\mathcal{F}_{t}$ is set to be the latest $\mathcal{G}_{i}$ of the active agent $i$ at round $t$. The detailed procedures of our FTDL are summarized in Algorithm 1 from the point of view of any single agent $i$. Despite of the simplicity, we prove FTDL enjoys the following regret bound, which demonstrates that it can exploit the strong convexity automatically.
@@
which actually is even one step ahead of FTL in (1). In the literature, it is well-known that such an ideal decision enjoys non-positive regret for any loss functions $f_{1}(\mathbf{x}),\dots,f_{T}(\mathbf{x})$ over $\mathcal{K}$.
@@
[…]
@@
By combining Lemma (1) with (3), it is easy to verify that
@@
$\|\mathbf{x}-\mathbf{x}^{*}\|^{2}_{2}\leq\frac{2}{\lambda}\left((f(\mathbf{x})-f(\mathbf{x}^{*})\right)$
@@
By combining the above inequality with the fact that any feedback from rounds $1,\dots,t-d$ must have been received by any agent $i$ at the end of round $t-1$, we have
@@
Remark.
First, from Theorem 1, our FTDL can exploit the strong convexity of functions to achieve an $O(d\log T)$ regret bound, which is much tighter than the $O(\sqrt{dT})$ regret bound achieved by the DDA algorithm (Hsieh et al., 2022). More interestingly, our FTDL does not require any prior information and assumption about the delay. Second, it is worth noting that the regret bound in Theorem 1 holds without Assumption 2, which implies that our FTDL can be directly applied to the unbounded set. The nice property actually inherits from the standard OCO with strongly convex functions (Hazan et al., 2007).
@@ 3.3 Approximate Follow-the-Delayed-Leader
Specifically, inspired by the approximate variant of FTL (Hazan et al., 2007), it is natural to apply our FTDL to the surrogate loss functions defined in (2), which adopts the following decision
@@
At first glance, to compute $\mathbf{x}_{t}$ in (6), these agents need to share both the gradient $\nabla f_{t}(\mathbf{x}_{t})$ and the decision $\mathbf{x}_{t}$ with others, which doubles the communication costs of the DDA algorithm (Hsieh et al., 2022). However, it is worth noting that $\mathbf{x}_{t}$ in (6) is equivalent to
@@
because the term $(\lambda/2)\|\mathbf{x}_{s}\|_{2}^{2}$ does not affect the minimizer of the optimization problem in (6). As a result, these agents only need to share one vector $\mathbf{z}_{t}=\nabla f_{t}(\mathbf{x}_{t})-\lambda\mathbf{x}_{t}$ with others, which enjoys the same communication costs as the DDA algorithm (Hsieh et al., 2022).
Furthermore, we notice that when $\mathcal{F}_{t}=\emptyset$, i.e., the active agent at round $t$ have not received any feedback yet, the decision $\mathbf{x}_{t}$ generated by (7) could be arbitrary in $\mathcal{K}$. To minimize the randomness, we introduce an additional regularization into (7) as below
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\mathbf{z}_{s},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}\right)+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}$
which unifies $\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}=\mathbf{0}$ for the case with $\mathcal{F}_{t}=\emptyset$ under Assumption 2. The detailed procedures of our approximate FTDL are summarized in Algorithm 2 from the point of view of any single agent $i$.
@@
Algorithm 2 Approximate FTDL (Agent $i$)
@@
3:     Asynchronously receive feedback $\mathbf{z}_{s}=\nabla f_{s}(\mathbf{x}_{s})-\lambda\mathbf{x}_{s}$ from round $s$
@@
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\mathbf{z}_{s},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|^{2}_{2}\right)+\frac{\lambda}{2}\|\mathbf{x}\|^{2}_{2}$
@@
$R_{T}\leq\frac{\lambda D^{2}}{2}+\frac{2d(G+2\lambda D)^{2}(\ln T+1)}{\lambda}.$
@@
For any $t=1,\dots,T$, let $\tilde{f}_{t}(\mathbf{x})=\langle\mathbf{z}_{t},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}$ and $F_{t}(\mathbf{x})=\sum_{s=1}^{t}\tilde{f}_{s}(\mathbf{x})+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}$. Similar to the proof of Theorem 1, for each round $t$, we can define an ideal decision as below
$\begin{split}\tilde{\mathbf{x}}_{t}=&\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s=1}^{t}\tilde{f}_{s}(\mathbf{x})+\frac{\lambda}{2}\|\mathbf{x}\|^{2}_{2}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}F_{t}(\mathbf{x}).\end{split}$ (8)
@@
Then, we notice that the first term in the right side of (9) can be bounded by analyzing the distance of our decision and the ideal decision. Specifically, under Assumptions 1 and 2, for any $t=1,\dots,T$, we have
$\begin{split}&\|\mathbf{z}_{t}\|_{2}=\|\nabla f_{t}(\mathbf{x}_{t})-\lambda\mathbf{x}_{t}\|_{2}\\ \leq&\|\nabla f_{t}(\mathbf{x}_{t})\|_{2}+\|\lambda\mathbf{x}_{t}\|_{2}\leq G+\lambda D.\end{split}$ (10)
Moreover, for any $t=1,\dots,T$, we can show that $\tilde{f}(\mathbf{x})$ is $(G+2\lambda D)$-Lipschitz, i.e., for any $\mathbf{x},\mathbf{y}\in\mathcal{K}$, it holds that
$\begin{split}\left|\tilde{f}_{t}(\mathbf{x})-\tilde{f}_{t}(\mathbf{y})\right|=&\left|\langle\mathbf{z}_{t},\mathbf{x}-\mathbf{y}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}-\frac{\lambda}{2}\|\mathbf{y}\|_{2}^{2}\right|\\ =&\left|\left\langle\mathbf{z}_{t}+\frac{\lambda}{2}(\mathbf{x}+\mathbf{y}),\mathbf{x}-\mathbf{y}\right\rangle\right|\\ \leq&\left\|\mathbf{z}_{t}+\frac{\lambda}{2}(\mathbf{x}+\mathbf{y})\right\|_{2}\|\mathbf{x}-\mathbf{y}\|_{2}\\ \leq&(G+2\lambda D)\|\mathbf{x}-\mathbf{y}\|_{2}\end{split}$ (11)
where the last inequality is due to (10) and Assumptions 1 and 2.
From (11), it is easy to verify that
$\begin{split}\sum_{t=1}^{T}\left(\tilde{f}_{t}(\mathbf{x}_{t})-\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})\right)\leq(G+2\lambda D)\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}.\end{split}$ (12)
Note that $F_{t}(\mathbf{x})$ is $((t+1)\lambda)$-strongly convex over $\mathcal{K}$. As a result, by combining Lemma 2 with the definition of $\tilde{\mathbf{x}}_{t}$ and $F_{t}(\mathbf{x})$, we have
@@
$\leq$ $\frac{2}{(t+1)\lambda}\left(F_{t}(\mathbf{x}_{t})-F_{t}(\tilde{\mathbf{x}}_{t})\right)$
$=$ $\frac{2}{(t+1)\lambda}\left(\sum_{s\in\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)+\frac{\lambda\left(\|\mathbf{x}_{t}\|_{2}^{2}-\|\tilde{\mathbf{x}}_{t}\|_{2}^{2}\right)}{2}\right)$
$+\frac{2}{(t+1)\lambda}\sum_{s\in[t]\setminus\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)$
$\leq$ $\frac{2}{(t+1)\lambda}\sum_{s\in[t]\setminus\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)$
$\leq$ $\frac{2\left(t-|\mathcal{F}_{t}|\right)(G+2\lambda D)}{(t+1)\lambda}\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}$
where the second inequality is due to the definition of $\mathbf{x}_{t}$, and the last inequality is due to (11).
By combining the above inequality with the fact any feedback from rounds $1,\dots,t-d$ must have been received by any agent $i$ at the end of round $t-1$, we have
$\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}\leq\frac{2d(G+2\lambda D)}{(t+1)\lambda}.$ (13)
Now, we only need to bound the last term in the right side of (9). Specifically, by applying Lemma 1 with the definition of $\tilde{\mathbf{x}}_{t}$, we have
$\begin{split}\left(\sum_{t=1}^{T}\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})+\frac{\lambda}{2}\|\tilde{\mathbf{x}}_{0}\|_{2}^{2}\right)-\left(\sum_{t=1}^{T}\tilde{f}_{t}(\mathbf{x}^{\ast})+\frac{\lambda}{2}\|{\mathbf{x}}^{\ast}\|_{2}^{2}\right)\leq 0\end{split}$
where $\tilde{\mathbf{x}}_{0}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}(\lambda/2)\|\mathbf{x}\|_{2}^{2}=\mathbf{0}$ under Assumption 1.
From the above inequality, it is not hard to verify that
$\begin{split}\sum_{t=1}^{T}\left(\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})-\tilde{f}_{t}(\mathbf{x}^{\ast})\right)\leq\frac{\lambda}{2}\|\mathbf{x}^{\ast}\|_{2}^{2}\leq\frac{\lambda D^{2}}{2}\end{split}$ (14)
@@
Finally, by combining (9) with (12), (13), and (14), we have
$\begin{split}R_{T}\leq&\frac{\lambda D^{2}}{2}+\sum_{t=1}^{T}\frac{2d(G+2\lambda D)^{2}}{(t+1)\lambda}\\ \leq&\frac{\lambda D^{2}}{2}+\frac{2d(G+2\lambda D)^{2}(\ln T+1)}{\lambda}.\end{split}$
Remark.
From Theorem 2, the approximate FTDL enjoys the same $O(d\log T)$ regret bound as FTDL, and also does not require any prior information or assumption about the delay. Moreover, we emphasize that although our Theorem 2 requires the boundedness of decision sets, i.e., Assumption 2, it is actually easy to extend our Algorithm 2 to further handle the unbounded set by using the fact that the optimal solution of a strongly convex function must lie in a ball centered at the origin (see Remark 2 of Wan et al. (2022b) for details).

### v2

@@ 1 Introduction
Motivated by these applications, Hsieh et al. (2022) extend the standard OCO into multi-agent asynchronous online optimization with delays. The setting involves a set of agents $\mathcal{M}=\{1,\ldots,M\}$, and only one agent becomes active in each round $t$ to make the decision $\mathbf{x}_{t}$. Due to the effect of delays and asynchronicities, each agent $i\in\mathcal{M}$ can only receive the feedback about $f_{t}(\cdot)$ at the end of round $t+d_{t,i}-1$, where $d_{t,i}\in\mathbb{Z}^{+}$ denotes an arbitrary delay. It is well-known that if only one agent exists, i.e., $M=1$, various algorithms have been proposed to achieve an optimal $O(\sqrt{dT})$ regret bound for convex functions and an optimal $O(d\log T)$ regret bound for strongly convex functions in the worst case (Weinberger and Ordentlich, 2002; Joulani et al., 2013; Quanrud and Khashabi, 2015; Joulani et al., 2016; Wan et al., 2022b), where $d=\max_{t,i}\{d_{t,i}\}$ denotes the maximum delay. However, compared with the single-agent case, multi-agent OCO with delays faces two new challenges: the non-monotonicity of feedback sequence and lack of synchronization, which make the time-decaying and even delay-dependent learning rates utilized in these single-agent delayed algorithms unavailable (Hsieh et al., 2022).
@@ 2.1 The Standard OCO
Since the pioneering work of Zinkevich (2003), there has been a growing research interest in developing and analyzing algorithms for OCO (Hazan et al., 2007; Abernethy et al., 2008; Duchi et al., 2011; Kingma and Ba, 2015; Wang et al., 2020; Wan and Zhang, 2021, 2022). Roughly speaking, these existing OCO algorithms can be divided into two categories. The first category includes online gradient descent (OGD) (Zinkevich, 2003) and its variants, which can be viewed as an extension from gradient descent in offline optimization (Boyd and Vandenberghe, 2004) into the online setting. The other category contains follow-the-leader (FTL) (Hannan, 1957) and its variants, where the so-called “leader” is the optimal decision for all historical loss functions. Note that both categories are able to achieve an $O(\sqrt{T})$ regret bound for convex functions, and an $O(\log T)$ regret bound for strongly convex functions (Zinkevich, 2003; Hazan et al., 2007; Hazan, 2016). Moreover, these $O(\sqrt{T})$ and $O(\log T)$ regret bounds have been proved to be optimal for convex and strongly convex functions, respectively Abernethy et al. (2008).
@@ 2.2 Multi-Agent OCO with Delays
where $\mathcal{F}_{t}$ is a set including the timestamps of all feedback received by the $t$-th active agent. Let $|\mathcal{F}_{t}|$ denotes the cardinality of $\mathcal{F}_{t}$ (i.e., the number of feedback received up to round $t$). They first prove that DDA can achieve the $O(\sqrt{dT})$ regret bound for convex functions, by setting
@@ 3.2 Follow-the-Delayed-Leader
Before introducing our algorithms, we first provide a formal definition for the available information held by each agent $i\in[M]$. To be precise, let $\mathcal{G}_{i}$ denote the set including the timestamps for all feedback received by each agent $i\in[M]$. In the beginning, the set should be initialized as $\mathcal{G}_{i}=\emptyset$. Then, due to the effect of delays and asynchronicity, each agent $i$ will receive the feedback from some previous round $s$, and update the set as $\mathcal{G}_{i}=\mathcal{G}_{i}\cup\{s\}$. Moreover, it will relay the received feedback if necessary. In this way, when an agent $i$ is activated at round $t$, it can only make the decision based on the feedback from rounds in the latest set $\mathcal{G}_{i}$ that may only be a subset of [t1][t-1].
Therefore, even if the functions are strongly convex, FTL in Eq. (1) cannot be directly utilized. To address this issue, we propose a delayed variant of FTL for multi-agent OCO with delays that is called follow-the-delayed-leader (FTDL) and utilizes the following decision
where $\mathcal{F}_{t}$ is set to be the latest $\mathcal{G}_{i}$ of the active agent $i$ at round $t$. Note that when $\mathcal{F}_{t}=\emptyset$, i.e., the active agent at round $t$ has not received any feedback yet, the decision $\mathbf{x}_{t}$ can be arbitrary in $\mathcal{K}$. The detailed procedures of our FTDL are summarized in Algorithm 1 from the point of view of any single agent $i$. Despite the simplicity, we prove FTDL enjoys the following regret bound, which demonstrates that it can exploit the strong convexity automatically.
Algorithm 1 Follow-the-Delayed-Leader (Agent ii)
@@
3:  Asynchronously receive feedback $f_{s}(\mathbf{x})$ of round $s$
4:  Update $\mathcal{G}_{i}=\mathcal{G}_{i}\cup\{s\}$
5:  if the agent becomes active then
6:   Set $\mathcal{F}_{t}=\mathcal{G}_{i}$
7:   Play $\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}f_{s}(\mathbf{x})$
8:  end if
9:end while
@@
which actually is even one step ahead of FTL in Eq. (1). In the literature, it is well-known that such an ideal decision enjoys non-positive regret for any loss functions $f_{1}(\mathbf{x}),\dots,f_{T}(\mathbf{x})$ over $\mathcal{K}$.
@@
$\ =$ $\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\tilde{\mathbf{x}}_{t})\right)+\sum_{t=1}^{T}\left(f_{t}(\tilde{\mathbf{x}}_{t})-f_{t}(\mathbf{x}^{\ast})\right)$
@@
By combining Lemma (1) with Eq. (3), it is easy to verify that
@@
$\|\mathbf{x}-\mathbf{x}^{*}\|^{2}_{2}\leq\frac{2}{\lambda}\left(f(\mathbf{x})-f(\mathbf{x}^{*})\right)$
@@
By combining the above inequality with the fact that any feedback from rounds $1,\dots,t-d$ must have been received by any agent $i$ at the end of round t1t-1, we have
@@
Remark 1.
First, from Theorem 1, our FTDL can exploit the strong convexity of functions to achieve an $O(d\log T)$ regret bound, which is much tighter than the $O(\sqrt{dT})$ regret bound achieved by the DDA algorithm (Hsieh et al., 2022). More interestingly, our FTDL does not require any prior information and assumption about the delay. Second, Assumption 2 is mainly introduced for making Assumption 1 rigorous. Actually, it is worth noting that the regret bound in Theorem 1 does not explicitly depend on the radius of $\mathcal{K}$ given by Assumption 2. The nice property actually inherits from the standard OCO with strongly convex functions (Hazan et al., 2007). Besides, we notice that this property naturally enables an extension of our algorithm to unconstrained domains, which will be detailed in Section 3.4.
@@ 3.3 Approximate Follow-the-Delayed-Leader
Specifically, inspired by the approximate variant of FTL (Hazan et al., 2007), it is natural to apply our FTDL to the surrogate loss functions defined in Eq. (2), which adopts the following decision
@@
At first glance, to compute $\mathbf{x}_{t}$ in Eq. (6), these agents need to share both the gradient $\nabla f_{t}(\mathbf{x}_{t})$ and the decision $\mathbf{x}_{t}$ with others, which doubles the communication costs of the DDA algorithm (Hsieh et al., 2022). However, it is worth noting that $\mathbf{x}_{t}$ in Eq. (6) is equivalent to
@@
because the term $(\lambda/2)\|\mathbf{x}_{s}\|_{2}^{2}$ does not affect the minimizer of the optimization problem in Eq. (6). As a result, these agents only need to share one vector $\mathbf{z}_{t}=\nabla f_{t}(\mathbf{x}_{t})-\lambda\mathbf{x}_{t}$ with others, which enjoys the same communication costs as the DDA algorithm (Hsieh et al., 2022). Therefore, the decision $\mathbf{x}_{t}$ can be computed as
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\mathbf{z}_{s},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}\right).$
The detailed procedures of our approximate FTDL are summarized in Algorithm 2 from the point of view of any single agent $i$.
@@
Algorithm 2 Approximate FTDL (Agent ii)
@@
3:  Asynchronously receive feedback $\mathbf{z}_{s}=\nabla f_{s}(\mathbf{x}_{s})-\lambda\mathbf{x}_{s}$ from the round $s$
@@
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\mathbf{z}_{s},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|^{2}_{2}\right)$
@@
$R_{T}\leq\frac{2d(G+2\lambda R)^{2}(\ln T+1)}{\lambda}.$
@@
For any $t=1,\dots,T$, let $\tilde{f}_{t}(\mathbf{x})=\langle\mathbf{z}_{t},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}$ and $F_{t}(\mathbf{x})=\sum_{s=1}^{t}\tilde{f}_{s}(\mathbf{x})$. Similar to the proof of Theorem 1, for each round $t$, we can define an ideal decision as below
$\begin{split}\tilde{\mathbf{x}}_{t}=&\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s=1}^{t}\tilde{f}_{s}(\mathbf{x})=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}F_{t}(\mathbf{x}).\end{split}$ (8)
@@
Then, we notice that the first term in the right side of Eq. (9) can be bounded by analyzing the distance of our decision and the ideal decision. Specifically, under Assumptions 1 and 2, for any $t=1,\dots,T$, we have
$\begin{split}&\|\mathbf{z}_{t}\|_{2}=\|\nabla f_{t}(\mathbf{x}_{t})-\lambda\mathbf{x}_{t}\|_{2}\\ \leq&\|\nabla f_{t}(\mathbf{x}_{t})\|_{2}+\|\lambda\mathbf{x}_{t}\|_{2}\leq G+\lambda R.\end{split}$ (10)
Moreover, for any $t=1,\dots,T$, we can show that $\tilde{f}(\mathbf{x})$ is $(G+2\lambda R)$-Lipschitz, i.e., for all $\mathbf{x}\in\mathcal{K}$, it holds that
$\begin{split}&\|\nabla\tilde{f}_{t}(\mathbf{x})\|_{2}\leq\|\mathbf{z}_{t}\|_{2}+\lambda\|\mathbf{x}\|_{2}\\ &\leq(G+\lambda R)+\lambda R=G+2\lambda R\end{split}$ (11)
where the second inequality is due to Eq. (10) and Assumption 2.
From Eq. (11), it is easy to verify that
$\begin{split}\sum_{t=1}^{T}\left(\tilde{f}_{t}(\mathbf{x}_{t})-\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})\right)\leq(G+2\lambda R)\sum_{t=1}^{T}\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}.\end{split}$ (12)
Note that $F_{t}(\mathbf{x})$ is $(t\lambda)$-strongly convex over $\mathcal{K}$. As a result, by combining Lemma 2 with the definition of $\tilde{\mathbf{x}}_{t}$ and $F_{t}(\mathbf{x})$, we have
@@
$=$ $\frac{2}{t\lambda}\sum_{s\in\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)$
$+\frac{2}{t\lambda}\sum_{s\in[t]\setminus\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)$
$\leq$ $\frac{2}{t\lambda}\sum_{s\in[t]\setminus\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)$
$\leq$ $\frac{2\left(t-|\mathcal{F}_{t}|\right)(G+2\lambda R)}{t\lambda}\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}$
where the second inequality is due to the definition of $\mathbf{x}_{t}$, and the last inequality is due to Eq. (11).
By combining the above inequality with the fact any feedback from rounds $1,\dots,t-d$ must have been received by any agent $i$ at the end of round t1t-1, we have
$\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}\leq\frac{2d(G+2\lambda R)}{t\lambda}.$ (13)
Now, we only need to bound the last term in the right side of Eq. (9). Specifically, by applying Lemma 1 with the definition of $\tilde{\mathbf{x}}_{t}$, it is not hard to verify that
$\begin{split}\sum_{t=1}^{T}\left(\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})-\tilde{f}_{t}(\mathbf{x}^{\ast})\right)\leq 0\end{split}$ (14)
@@
Finally, by combining Eq. (9) with Eq. (12), Eq. (13), and Eq. (14), we have
$\begin{split}R_{T}\leq&\sum_{t=1}^{T}\frac{2d(G+2\lambda R)^{2}}{t\lambda}\\ \leq&\frac{2d(G+2\lambda R)^{2}(\ln T+1)}{\lambda}.\end{split}$
Remark 2.
From Theorem 2, the approximate FTDL enjoys the same $O(d\log T)$ regret bound as FTDL, and also does not require any prior information and assumption about the delay. However, we want to clarify that our A-FTDL actually also relies on the delay via the set $\mathcal{F}_{t}$. Although such a reliance is not easily noticeable, it actually plays a role similar to a delay-dependent step-size used in DDA (Hsieh et al., 2022). Moreover, we emphasize that although Theorem 2 requires the boundedness of decision sets, i.e., Assumption 2, it is actually easy to extend our Algorithm 2 to further handle the unbounded set by using the fact that the optimal solution of a strongly convex function must lie in a ball centered at the origin, as detailed in Section 3.4.
Remark 3.
We want to clarify that in A-FTDL the argmin subproblem that we need to solve is equal to a projection operation onto the convex set $\mathcal{K}$, i.e.,
$\mathbf{x}_{t}$ $=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\mathbf{z}_{s},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|^{2}_{2}\right)$
$=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\left\{\frac{\lambda}{2}\left\|\mathbf{x}+\frac{\sum_{s\in\mathcal{F}_{t}}\mathbf{z}_{s}}{\lambda}\right\|^{2}_{2}-\frac{1}{2\lambda}\left\|\sum_{s\in\mathcal{F}_{t}}\mathbf{z}_{s}\right\|^{2}_{2}\right\}$
$=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\left\|\mathbf{x}+\frac{\sum_{s\in\mathcal{F}_{t}}\mathbf{z}_{s}}{\lambda}\right\|^{2}_{2}\$
This computational cost is also required by the previous DDA algorithm (Hsieh et al., 2022), and many other online optimization algorithms.